OpenStudy (anonymous):

If g(x) = 3x - 5, then g^-1(x)=

OpenStudy (anonymous):

g^-1(x)=(x+5)/3

OpenStudy (anonymous):

To find \[g^{-1}\] We write : \[g(y)=x\] and we write y like this : \[3y-5=x\\ 3y=x+5\\ y=\frac{x+5}{3} \] So : \[g^{-1}(x)=\frac{x+5}{3}\]